# FollowSpot: Enhancing Wireless Communications via Movable Ceiling-Mounted Metasurfaces

### v1

@@ III-B Global Solution in Polynomial Time
We have shown that Algorithm 1 can recover the corresponding $\bm{X}^{\star}$ given a particular $\mu$. Since the optimal $\mu$ must reside in one of the different $\xi$-arcs, the solution produced by Algorithm 1 must be optimal for problem (13). The above theorem immediately follows by the equivalence between problem (5) and problem (13).
III-C A Specific Example
TABLE II: Reflected channels considered in the toy example.
$m=1$ $m=2$
$\ell=1$ $(0.6+0.6\mathrm{j})\times 10^{-6}$ $(-0.5+0.2\mathrm{j})\times 10^{-6}$
$\ell=2$ $(-1.1+0.8\mathrm{j})\times 10^{-6}$ $(0.9+1.3\mathrm{j})\times 10^{-6}$
$\ell=3$ $(1.2-1.4\mathrm{j})\times 10^{-6}$ $(-1.3+0.7\mathrm{j})\times 10^{-6}$
We now illustrate how Algorithm 1 works through a toy example. Let $M=2$ and $L=3$. The direct channel $h_{0}=2.0\times 10^{-6}$; the reflected channels are shown in Table II. For $m=1$, we consider solving equation (15) for (,)$(\ell,\ell^{\prime})\in\{(1,2),(1,3),(2,3)\}$. By (16) and (17), we obtain six solutions to (15):
$\{e^{\mathrm{j}0.61\pi},e^{\mathrm{j}1.61\pi},e^{\mathrm{j}0.73\pi},e^{\mathrm{j}1.73\pi},e^{\mathrm{j}0.88\pi},e^{\mathrm{j}1.88\pi}\}.$ (18)
Likewise, for $m=2$, we obtain another set of six solutions:
$\{e^{\mathrm{j}0.28\pi},e^{\mathrm{j}1.28\pi},e^{\mathrm{j}0.40\pi},e^{\mathrm{j}1.40\pi},e^{\mathrm{j}0.68\pi},e^{\mathrm{j}1.68\pi}\}.$ (19)
Putting together the above solutions and sorting them out, we obtain the ordered sequence of $\xi_{\ell}$:
$\begin{matrix}\xi_{1}=e^{\mathrm{j}0.28\pi},&\xi_{2}=e^{\mathrm{j}0.40\pi},&\xi_{3}=e^{\mathrm{j}0.61\pi},&\xi_{4}=e^{\mathrm{j}0.68\pi},\\ \xi_{5}=e^{\mathrm{j}0.73\pi},&\xi_{6}=e^{\mathrm{j}0.88\pi},&\xi_{7}=e^{\mathrm{j}1.28\pi},&\xi_{8}=e^{\mathrm{j}1.40\pi},\\ \xi_{9}=e^{\mathrm{j}1.61\pi},&\xi_{10}=e^{\mathrm{j}1.68\pi},&\xi_{11}=e^{\mathrm{j}1.73\pi},&\xi_{12}=e^{\mathrm{j}1.88\pi}.\end{matrix}$
The above points divide the unit circle into $12$ $\xi$-arcs as shown in Fig. 4. Next, we place $\mu$ at the midpoint of each $\xi$-arc, compute the corresponding $\bm{X}^{\star}$ according to (11), and also compute the resulting value of $f(\bm{X}^{\star})$. These computation results are summarized in Table III. Since the $12$th $\xi$-arc yields the largest value of $f(\bm{X}^{\star})$, we conclude that
$\bm{X}^{\star}=\begin{bmatrix}0&0&1\\ 0&1&0\end{bmatrix}$
is the optimal solution to problem (5).
TABLE III: Computation results of the toy example in Section III-C.
$\ell$ 1 2 3 4 5 6 7 8 9 10 11 12
$\mu$ $e^{\mathrm{j}0.34\pi}$ $e^{\mathrm{j}0.51\pi}$ $e^{\mathrm{j}0.65\pi}$ $e^{\mathrm{j}0.71\pi}$ $e^{\mathrm{j}0.81\pi}$ $e^{\mathrm{j}1.08\pi}$ $e^{\mathrm{j}1.34\pi}$ $e^{\mathrm{j}1.51\pi}$ $e^{\mathrm{j}1.65\pi}$ $e^{\mathrm{j}1.71\pi}$ $e^{\mathrm{j}1.81\pi}$ $e^{\mathrm{j}0.08\pi}$
$\bm{X}^{\star}$ $\left[\begin{smallmatrix}0&0&1\\ 1&0&0\end{smallmatrix}\right]$ $\left[\begin{smallmatrix}0&0&1\\ 0&0&1\end{smallmatrix}\right]$ $\left[\begin{smallmatrix}0&1&0\\ 0&0&1\end{smallmatrix}\right]$ $\left[\begin{smallmatrix}0&1&0\\ 0&1&0\end{smallmatrix}\right]$ $\left[\begin{smallmatrix}1&0&0\\ 0&1&0\end{smallmatrix}\right]$ $\left[\begin{smallmatrix}0&0&1\\ 0&1&0\end{smallmatrix}\right]$
$f$ $2.9\times 10^{-6}$ $2.0\times 10^{-6}$ $1.6\times 10^{-6}$ $2.8\times 10^{-6}$ $3.9\times 10^{-6}$ $4.1\times 10^{-6}$
Figure 4: In the toy example with $M=2$ and $L=3$, we have 12 solutions $\{\xi_{\ell}\}$ to equation (15) for $m\in\{1,2\}$ and (,)$(\ell,\ell^{\prime})\in\{(1,2),(1,3),(2,3)\}$. These solutions partition the unit circle into $12$ $\xi$-arcs.
@@ IV Extension to Multi-Receiver Case
The goal of this section is to extend Algorithm 1 to multiple actuators. Following the industrial communication protocol called the Occupy CoW [45], we assume that the controller broadcasts a common message to $U\geq 2$ actuators. Denote by $h_{0,u}$ the direct channel from the controller to the $u$th actuator, and $h_{m,\ell,u}$ the reflected channel induced by MTS $m$ for the $u$th actuator when the MTS is placed at position $\ell$ in its cell, where $m\in\mathcal{M}$, $\ell\in\mathcal{L}_{m}$, and $u=1,\ldots,U$. The resulting SNR of actuator $u$ is then given by
@@
We propose a natural extension of Algorithm 1 based on majority voting [46, 47] to solve problem (21) sub-optimally. First, run Algorithm 1 for each user $u$ separately. Let $\bm{X}^{(u)}$ be the solution of Algorithm 1 when actuator $u$ is the target user in consideration. As before, let $\bm{x}^{(u)}_{m}$ be the $m$th row vector of $\bm{X}^{(u)}$, and let $x^{(u)}_{m,\ell}$ be the $\ell$th entry of $\bm{x}^{(u)}_{m}$. Clearly, $\bm{X}^{(u)}$ obtained for different actuators are usually not equal. We now adopt a majority voting strategy to reach a consensus. Specifically, for each $\bm{x}_{m}$ (which represents the placement of MTS $m$), we collect the decisions of $\bm{x}^{(u)}_{m}$ from the different $u$, and place the MTS at the position that receives the most votes:
@@ V Numerical Results
We now numerically evaluate the performance of the proposed algorithms. Consider a $100\times 20$ ceiling area in meters, which is divided into $M=M_{x}\times M_{y}$ equal rectangular cells as shown in Fig. 2. Each cell provides $L$ candidate positions for its MTS. The ceiling height is $5$ meters. The actuator(s) are randomly distributed on the ground. We set $M=6\times 5=30$, $L=6$, and $N=100$ by default.
@@
$h=\sqrt{\gamma}\left(\sqrt{\frac{\delta}{1+\delta}}\overline{h}+\sqrt{\frac{1}{1+\delta}}\widetilde{h}\right),$ (24)
where $10<\gamma<$ 1 is the pathloss factor, $\delta>0$ is the Rician factor, $\overline{h}\in\mathbb{C}$ is the normalized line-of-sight (LoS) component, and the fading component $\widetilde{h}$ is a random variable drawn from the complex Gaussian distribution $\mathcal{CN}(0,1)$. The parameter of the pathloss follow [48, 49, 50], which is generated according to
@@
Figure 7: SNR boost vs. the number of meta-atoms $N$.
@@
  • Fix-Placement: Fix the positions of $M$ MTSs.
@@
SNR boost=SNR with MTSs deployedSNR without $\text{SNR boost}=\frac{\text{SNR with MTSs deployed}}{\text{SNR without MTS}},$
@@
Fig. 8 further shows the SNR boost achieved by different algorithms versus the number of candidate positions $L$ for each MTS. It can be seen that increasing $L$ can also lead to performance improvement of different algorithms (except the Fix-Placement method). Observe that, as $L$ increases, the performance of Algorithm 1 and the CMP method first improves rapidly and then gradually stabilizes. This trend indicates that the performance gain from further increasing the number of candidate positions is limited. Observe also that, when $L=2$, the performance of the RMP is much better than that of the CMP. But when $L\geq 4$, the performance of RMP is worse than CMP. We also observe that the proposed Algorithm 1 always achieves the best performance regardless of the values of $L$.
@@
Finally, we compare the running time of different algorithms in both the single-receiver and the multi-receiver cases as shown in Table IV. It can be seen that the runtimes of the proposed Algorithm 1 and Algorithm 2 are slightly higher than that of the CMP method but lower than that of the RMP method. We also compared the runtime of the Algorithm 1 with exhaustive search on a toy model with $M=10$, $L=6$, and $U=1$. Our Algorithm 1 completes in $0.0023$ seconds, whereas exhaustive search takes $13.0053$ seconds. These results suggest that the proposed algorithm has low computational complexity, so it is well-suited for implementation on resource-constrained devices.
TABLE IV: Runtimes (in Second) of different algorithms

### v2

@@ III-B Global Solution in Polynomial Time
We have shown that Algorithm 1 can recover the corresponding $\bm{X}^{\star}$ given a particular $\mu$. Since the optimal $\mu$ must reside in one of the different $\xi$-sectors, the solution produced by Algorithm 1 must be optimal for problem (13). The above theorem immediately follows by the equivalence between problem (5) and problem (13).
@@ IV Extension to Multi-Receiver Case
The goal of this section is to extend Algorithm 1 to multiple actuators. We assume that the controller broadcasts a common message to $U\geq 2$ actuators. If one wishes to deliver separate messages to multiple actuators, we propose using the so-called “Occupy CoW” scheme [41, 24, 3, 8] (which was proposed exactly for the industrial communications) to somehow convert the separate-message broadcast to the common-message broadcast. Denote by $h_{0,u}$ the direct channel from the controller to the $u$th actuator, and $h_{m,\ell,u}$ the reflected channel induced by MTS $m$ for the $u$th actuator when the MTS is placed at position $\ell$ in its zone, where $\in\mathcal{M}, m\ell\in\mathcal{L}_{m}$, and $u=1,\ldots,U$. The resulting SNR of actuator $u$ is then given by
@@
[…]
@@
Remark 2
The proposed MTS placement method can be readily combined with the existing algorithms for phase shifting and power transmission in an alternating fashion. Specifically, after the positions of MTSs have been determined, we can refine the CSI measure and then perform any existing algorithms for phase shifting and power transmission.
@@ V Numerical Results
We now numerically evaluate the performance of the proposed algorithms. Consider a $100\times 20$ ceiling in meters, which is divided into $M=M_{x}\times M_{y}$ equal rectangular zones as shown in Fig. 2. Each zone provides $L$ candidate positions for its MTS. The ceiling height is $5$ meters. The actuator(s) are randomly distributed on the ground. We set $M=6\times 5=30$, $L=6$, and $N=100$ by default.
@@
[…]
where $0<\gamma<1$ is the pathloss factor, $\delta>0$ is the Rician factor, $\overline{h}\in\mathbb{C}$ is the normalized line-of-sight (LoS) component, and the fading component $\widetilde{h}$ is a random variable drawn from the complex Gaussian distribution $\mathcal{CN}(0,1)$. The parameter of the pathloss follow [38, 50, 20], which is generated according to
@@
Figure 7: SNR boost vs. the number of meta-atoms NN.
@@
  • Fix-Placement: Fix the randomly assigned positions of the $M$ MTSs.
@@
$\text{SNR boost}=\frac{\text{SNR with MTSs deployed}}{\text{SNR without MTS}},$
@@
Fig. 8 further compares the performance of the proposed method across different grid resolutions, as indicated by varying values of $L$. It can be seen that increasing $L$ can also lead to performance improvement of different algorithms (except the Fix-Placement method). Observe that, as $L$ increases, the performance of Algorithm 1 and the CMP method first improves rapidly and then gradually stabilizes. This trend indicates that the performance gain from further increasing the number of candidate positions is limited. Observe also that, when $L=2$, the performance of the RMP is much better than that of the CMP. But when $L\geq 4$, the performance of RMP is worse than CMP. We also observe that the proposed Algorithm 1 always achieves the best performance regardless of the values of $L$.
@@
Finally, we compare the running time of different algorithms in both the single-receiver and the multi-receiver cases as shown in Table II. It can be seen that the runtimes of the proposed Algorithm 1 and Algorithm 2 are slightly higher than that of the CMP method but lower than that of the RMP method. We also compared the runtime of the Algorithm 1 with exhaustive search on a toy model with $M=10$, $L=6$, and $U=1$. Our Algorithm 1 completes in $0.0023$ seconds, whereas exhaustive search takes $13.0053$ seconds. These results suggest that the proposed algorithm has low computational complexity, so it is well-suited for implementation on resource-constrained devices.
TABLE II: Runtimes (in Second) of different algorithms